OpenStudy (anonymous):

As part of your job as a quality-control inspector in a factory you can check 16 parts in 3 minutes. How long will it take you to check 80 parts? a. 15 minutes b. 26.6 minutes c. 5 minutes d. 74 minutes

OpenStudy (holly00d1248):

A

OpenStudy (krazykricket):

@Holly00d1248 please show her how you got that answer and help her through the process. If you don't she does get an answer and a good grade if it right, but she learns nothing from it.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yeah i hate words problems and i never get them right

OpenStudy (holly00d1248):

do 80/16=5*3=15

OpenStudy (holly00d1248):

Does this confuse you or help @rocky1718

OpenStudy (krazykricket):

@rocky1718 It's probably cause normally you see a short equation and your like oh this is easy, and then you come upon this paragraph and your mind doesn't concentrate and it's to busy trying to read it more then once while you try and solve the problem. All you have to do is relax and go slow.

OpenStudy (krazykricket):

Thanks @Holly00d1248 :) I know some people come on here just for answers and expect to learn things from it, but they don't.

OpenStudy (holly00d1248):

Ya i'm with @KrazyKricket i'm lazy when it comes to reading and sure @KrazyKricket

OpenStudy (anonymous):

ok so it will be 16/3 = 5.333333 but is not it if i divided

OpenStudy (holly00d1248):

no 80/16

OpenStudy (anonymous):

ok i see what you did

OpenStudy (holly00d1248):

OK

OpenStudy (anonymous):

80/16 is 5 multipli by 3 is 15

OpenStudy (krazykricket):

@rocky1718 If I was able to give more then one medal, I would give you one to. Since @Holly00d1248 was the helper I had to give it to him for his help and time.

OpenStudy (anonymous):

is ok thx btw

OpenStudy (krazykricket):

Your very Welcome :) If you ever need help, and no one comes within 2 minutes. Tag me in it, I may know it I may not.

OpenStudy (anonymous):

ok

OpenStudy (holly00d1248):

I gave you a medal @rocky1718 and i'm always open to help to tag me whenever you feel like it

OpenStudy (holly00d1248):

So need anymore help